OpenStudy (anonymous):

y^2 is directly proportional to x^3. If y is doubles find the percentage increase in x

OpenStudy (anonymous):

\[y^2=kx^3\] for some real \(k\). When \(y\) doubles: \[(2y)^2=4y^2=4kx^3=k\left(4^{1/3}x\right)^3\approx k(1.587x)^3\] The percentage increase is approximately 59%, because \(1.59x=x+\color{red}{.59x}\).